OpenStudy (anonymous):

what is 3 and 1/4 + 2 and 1/8

OpenStudy (anonymous):

5.375 or 3/48 in fraction form :)

OpenStudy (skullpatrol):

@sunnyshores why are you giving only an answer?

OpenStudy (skullpatrol):

$$\Huge 3 \frac{1}{4} + 2\frac{1}{8}$$ $$ \Huge= 3 +\frac{1}{4} + 2+\frac{1}{8}$$

OpenStudy (skullpatrol):

Recall a + b = b + a